OpenStudy (anonymous):

find the values of a & b that would make the quadrilateral a parallelogram (pic shown)

OpenStudy (anonymous):



OpenStudy (binary3i):

in an parallelogram opposite angles are equal 5b+6=8a-10 and adjacent angles adds up to 90 (4a-8)+(8a-10)=90 now you have the two equations to solve for a and b

OpenStudy (anonymous):

ohhh okay. i knew those two had to equal each other.. my answer just never made sense! thanks !

OpenStudy (binary3i):

so you got it?

OpenStudy (jdoe0001):

(4a-8)+(8a-10)=180?

OpenStudy (anonymous):

the answer i got was different than the back of the book :/

OpenStudy (jdoe0001):

try with 5b+6=8a-10 (4a-8)+(8a-10)=180

OpenStudy (anonymous):

now i got the right answer! thanks!

OpenStudy (binary3i):

oh yeah im so sorry i made a mistake they should add up to 180 and not 90.

OpenStudy (anonymous):

its fine :) but thanks again